OpenStudy (canada907cat):

A commuter is driving along a highway on which the speed limit is 45 miles per hour when he unknowingly runs into a speed trap involving two police officers. The first officer is positioned at mile marker 145 and clocks the commuter’s car at 45 miles per hour. Five minutes later, a second police officer at mile marker 149 clocks the car at 40 miles per hour. Can the commuter be charged with a speeding violation?

OpenStudy (canada907cat):

This is what I have:( 4miles/5minutes) * (60minutes/1 hour)=48 mi/hr. If this correct?

OpenStudy (canada907cat):

Total distance: 149-145=4 miles Total time= 5 min= 5/60hour=1/12 hour average speed= 4/(1/12)=4*12= 48 miles per hour

OpenStudy (canada907cat):

*is

OpenStudy (canada907cat):

@zepdrix